OpenStudy (anonymous):

Simplify the expression: 3 log 3 (6x – 4)

OpenStudy (anonymous):

3 log 3 (6x – 4) = 18 x log(3)-12 log(3) = 6 (3 x-2) log(3)

OpenStudy (anonymous):

Is it a \[3\log_{3}(6x-4) \]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Is the base of the log given? Or just an empty space at where the base should be?

OpenStudy (anonymous):

just an empty space. OK how about solving for if it's log to base 3 as well? :)

OpenStudy (anonymous):

If it's empty then it means a \[log_{10}\] So I'm thinking if the simplification can be \[lg 27+ lg(6x-4)^{3}\] Note: \[lg = log_{10}\]

OpenStudy (anonymous):

OK how about solving for if it's log to base 3 as well? :)

OpenStudy (anonymous):

Then it'll involve change of base... Great... Suddenly forgot how to solve that... I know but I'm not sure if it's correct

OpenStudy (anonymous):

any one to help me solve \[3\log _{3}(6x-4)\]?